Tshego won R165000 and decided to deposit 65% of this amount in an account earning 8,25% interest per year, compounded every four months. What is the accumulated amount after five years?

1
Expert's answer
2022-02-15T13:50:40-0500

SOLUTION:

Total money = R165000

65÷100 × 165000

= 107250

Money deposited = R107250

Rate = 8.25% per annum (p. a)

(A year has three periods of four months)

Rate per four months = 8.25÷3

= 2.75%

period = 5 (years) × 3

= 15

A=P(1+(r÷100))n

= 107250 × (1+(2.75÷100))15

= 107250 × (1.0275)15

= 107250 × 1.5022

Amount = R161110.95